# Mario Shannon

### v1

@@ Abstract.
For an Anosov flow on a closed 3-manifold, we are interested in the problem of whether or not making non-trivial Fried surgeries along a finite set of periodic orbits can produce a flow equivalent to itself. We show that for some suspension Anosov flows, there exist infinitely many pairs of periodic orbits satisfying this property.
@@ 1. Introduction
Given an Anosov flow on a closed 3-manifold and a periodic orbit, there is an operation called Fried surgery that allows to construct a new Anosov flow on a new 3-manifold, by making a modification of the initial flow around the chosen periodic orbit. This modification is essentially a Dehn surgery along the periodic orbit, but adapted to the pairs (Anosov flow, 3-mfld) in such a way that the resulting manifold is also equipped with an Anosov flow.
As a consequence, this operation allows to define a graph structure on the set of (orbital equivalence classes of) Anosov flows, where edges connecting two pairs (Anosov flow, 3-mfld) represent Fried surgeries that allow to obtain one flow from the other.
The purpose of this note is to show that, at some vertices of the graph corresponding to suspension Anosov flows, there exists infinitely many closed paths of length $=2$. This replies to a statement made in [7], where the authors claim that there should exist at most finitely many loops of this kind.
@@ 1.1. Fried Surgery on Anosov flows
A non-singular flow $\phi_{t}:M\to M$ on a closed Riemannian 3-manifold is Anosov if it is generated by a smooth vector field $X$ and the action $D\phi_{t}:TM\to TM$ on the tangent bundle preserves a uniformly hyperbolic splitting $TM=E^{s}\oplus E^{c}\oplus E^{u}$, where $E^{s}$ (resp. $E^{u}$) is a line-bundle exponentially contracting (resp. expanding) under the $D\phi_{t}$-action, and $E^{c}$ is a line-bundle colinear with $X$.
We refer to [13] for general properties about dynamics preserving a uniformly hyperbolic splitting. Among many interesting properties, Anosov flows always have infinitely many periodic trajectories, all of them saddle type hyperbolic. The Stable Manifold Theorem asserts that the plane bundles $E^{s}\oplus E^{c}$ and $E^{c}\oplus E^{u}$ are tangent to a pair of transverse codimension $=1$ foliations $\mathcal{F}^{cs}$ and $\mathcal{F}^{cu}$ on $M$, that intersects along the $\phi_{t}$-orbits. We call them center-stable and center-unstable foliations. Every leaf of these foliations is homeomorphic to either a cylinder or a Mobius band (in case the leaf contains a periodic orbit), or to a plane.
Let $\gamma$ be a periodic orbit of an Anosov flow $\phi_{t}:M\to M$ and assume $\mathcal{F}^{cs}(\gamma)$, $\mathcal{F}^{cu}(\gamma)$ are cylinders. In [12] Fried introduced a surgery operation, represented in figure 1, that consists in:
@@
    Choose a blow-up $(\phi_{t}^{*},M^{*})$ of the flow $(\phi_{t},M)$ along $\gamma$, whose restriction to the boundary $\phi_{t}^{*}:\partial M^{*}\to\partial M^{*}$ is Morse-Smale.
    A blow-up of the pair $(\phi_{t},M)$ along $\gamma$ is a 3-manifold $M^{*}$ with boundary homeomorphic to a torus, equipped with a flow $\phi_{t}^{*}:M^{*}\to M^{*}$ and a map $\pi:M^{*}\to M$ such that: (i)- $\pi:\partial M^{*}\to\gamma$ is a $\mathbb{S}^{1}$-fibration; (ii)- $\pi:M^{*}\setminus\partial M^{*}\to M\setminus\gamma$ is a homeomorphism; and (iii)- $\pi\circ\phi_{t}^{*}=\phi_{t}\circ\pi$, $\forall\ t\in\mathbb{R}$. There always exists a blow-up where the flow on the boundary is a Morse-Smale flow with two attracting and two repelling periodic orbits.
@@
    For this, choose a foliation $\mathcal{C}$ by simple closed curves on $\partial M^{*}$ such that each leaf is transverse to the flow $\phi_{t}^{*}:\partial M^{*}\to\partial M^{*}$ and intersects once each periodic orbit in the boundary. Up to a time reparametrization near the boundary we may assume that $\phi_{t}^{*}$ preserves this foliation. Then, the quotient $\sigma:M^{*}\to N\coloneqq M^{*}/\mathcal{C}$ is a $3$-manifold where $\delta\coloneqq\partial M^{*}/\mathcal{C}$ is a simple closed curve, and there is an induced flow $\psi_{t}:N\to N$ having $\delta$ as a saddle type periodic orbit.
Figure 1. Fried surgery $(\phi_{t},M)\xrightarrow[]{(\gamma,1)}(\psi_{t},N)$.
Topologically, this operation is an integral Dehn surgery on $M$ along the simple closed curve $\gamma$. Thus the homeomorphism type of $N$ is determined by the homology class in $\partial M^{*}$ of the leaves $C$ of the chosen foliation $\mathcal{C}$. In the basis of $H_{1}(\partial M^{*})$ given by the class $a$ of an oriented $\mathbb{S}^{1}$-fiber (meridian) and the class $b$ of an oriented closed orbit (longitude) we have that $[C]=a+m\cdot b$ for some $m\in\mathbb{Z}$.
This surgery has the additional property that the resulting manifold is equipped with a flow $\psi_{t}:N\to N$, that is actually equivalent to $(\phi_{t},M\setminus\gamma)$ in the complement of $\delta$. It is possible to check that the orbit $\delta$ is saddle type, and that the flow $\psi_{t}$ preserves a pair of globally defined center-stable and center-unstable foliations on $N$, induced from those corresponding to $\phi_{t}:M\to M$. However, this flow is not an Anosov flow (i.e. does not preserves a hyperbolic splitting) in a canonical way (see [17] for a deeper discussion). It is what is called a topological Anosov flow.
@@
We say that the previous operation is an integral Fried surgery on the flow $(\phi_{t},M)$ along the orbit $\gamma$ with slope $=m$. We denote it by
$(\psi_{t},M)=\textsl{Surg}(\phi_{t},M;(\gamma,m)).$
We denote by $\mathcal{P}:\textrm{Per}(\phi_{t},M)\to\textrm{Per}(\psi_{t},M)$ the natural map that associates periodic orbits before and after surgery.
The previous surgery can be also performed if $\mathcal{F}^{cs}(\gamma)$, $\mathcal{F}^{cu}(\gamma)$ are Mobius bands. However, in the present note we will only work with Anosov flows having orientable center-stable/unstable foliations.
@@ 1.2. Graph of Anosov flows
We say that two flows are orbitally equivalent if there exists a homeomorphism between their supporting manifolds that sends the orbit foliation of one flow to the orbit foliation of the other, preserving the orientation of the orbits. It is direct to check that Fried surgeries are well-defined up to orbital equivalence in the family of (topological) Anosov flows on closed 3-manifolds. This allows to define a graph structure as follows:
@@
Define the graph of Anosov flows to be the 1-complex $\mathbf{G}$ having:
  • Vertices $\mathbf{G}^{0}$: orbital equivalence classes of Anosov flows $(\phi_{t},M)$;
  • Edges $\mathbf{G}^{1}$: Fried surgeries $(\phi_{t},M)\xrightarrow[]{(\gamma,m)}(\psi_{t},N)$, $m\in\mathbb{Z}$ (non-oriented).
Anosov flows can be transitive or not, their $cs/cu$-foliations can be orientable or not, and their supporting 3-manifold can be orientable or not. Since Fried surgeries preserve each of these properties, the graph $\mathbf{G}$ has many connected components. One fundamental open problem is the following:
Problem 4 (Fried-Ghys).
Let $\mathbf{G}_{\circ}$ be the subgraph of $\mathbf{G}$ whose vertices represent Anosov flows on orientable closed 3-manifolds, dynamically transitive and having orientable $cs/cu$-foliations. Is it connected ?
One way to investigate the topology of $\mathbf{G}$ consists in fixing one Anosov flow, a set of $n>0$ periodic orbits, and look at all possible surgeries along those curves. This approach is studied by Bonatti and Iakovoglou in [7], for the case where the Anosov flow is a suspension (cf. section 2.1) and $n=2$.
Theorem 5 (Bonatti-Iakovouglu).
Let $(\phi_{t}^{A},M_{A})$ be the suspension flow generated by a hyperbolic matrix $A\in\mathrm{SL}(2,\mathbb{Z})$ with $\textsc{tr}(A)\geq 3$. Then,
@@
    For every pair of different periodic orbits $\Gamma=\{\gamma_{1},\gamma_{2}\}$ the set of $(m_{1},m_{2})$ in $\mathbb{Z}^{2}$ such that $\textsl{Surg}(\phi_{t}^{A},M_{A};(\gamma_{1},m_{1}),(\gamma_{2},m_{2}))$ is equivalent to a suspension Anosov flow, is finite. (Proposition 9.1 from [7])
@@
    There exist pairs $\Gamma=\{\gamma_{1},\gamma_{2}\}$ such that $\textsl{Surg}(\phi_{t}^{A},M_{A};(\gamma_{1},m_{1}),(\gamma_{2},m_{2}))$ is equivalent to a suspension Anosov flow if and only if $(m_{1},m_{2})=(0,0)$. (Theorem 5 from [7])
@@
Along the discussion in section 1.3 of [7] the authors claim that, for every fixed hyperbolic matrix $A\in\mathrm{SL}(2,\mathbb{Z})$, it seems plausible the existence of at most finitely many 4-tuples $(\gamma_{1},m_{1},\gamma_{2},m_{2})$ with $(m_{1},m_{2})\neq(0,0)$ such that the surgery
$\textsl{Surg}(\phi_{t}^{A},M_{A};(\gamma_{1},m_{1}),(\gamma_{2},m_{2}))$
produces a suspension Anosov flow.
In this note we prove that this is not the case, by showing the existence of infinitely many different paths of length $=2$ in $\mathbf{G}_{\circ}$ starting at $(\phi_{t}^{A},M_{A})$ and ending in another suspension.
1.3. $\infty$-many loops of length $=2$ based at a suspension.

### v2

# Infinitely many closed paths in the graph of Anosov flows

@@ Abstract.
Given an Anosov flow on a closed 3-manifold, we are interested in the problem of whether or not making non-trivial Fried surgeries along a finite set of periodic orbits can produce a flow equivalent to itself. We show that for some suspension Anosov flows, there exist infinitely many pairs of periodic orbits satisfying this property.
Key words and phrases:
Anosov flows, Dehn surgery, partially hyperbolic dynamics
2020 Mathematics Subject Classification:
37D20, 57M50
@@ 1. Introduction
Given an Anosov flow on a closed 3-manifold and a periodic orbit, there is an operation called Fried surgery that allows to construct a new Anosov flow on a new 3-manifold, by making a modification of the initial flow around the chosen periodic orbit. This modification is essentially a Dehn surgery along the periodic orbit, but adapted to the pairs (Anosov flow, 3-manifold) in such a way that the resulting manifold is also equipped with an Anosov flow.
As a consequence, this operation allows to define a graph structure on the set of (orbital equivalence classes of) Anosov flows, where edges connecting two pairs (Anosov flow, 3-manifold) represent Fried surgeries that allow to obtain one flow from the other.
The purpose of this note is to show that, at some vertices of the graph corresponding to suspension Anosov flows, there exist infinitely many closed paths of length 2. This replies to a statement made in [7] by Bonatti and Iakovoglou, where those authors claim that there should exist at most finitely many loops of this kind.
@@ 1.1. Fried Surgery on Anosov flows
A non-singular flow $\phi=\{\phi_{t}\}_{t\in\mathbb{R}}$ on a closed Riemannian 3-manifold $M$ is Anosov if it is generated by a smooth vector field $X$ and the action $D\phi_{t}:TM\to TM$ on the tangent bundle preserves a uniformly hyperbolic splitting $TM=E^{\mathrm{s}}\oplus E^{\mathrm{c}}\oplus E^{\mathrm{u}}$, where the $E^{\mathrm{s}}$ (resp. $E^{\mathrm{u}}$) line-bundle is exponentially contracted (resp. expanded) under the $D\phi_{t}$-action for $t\geq 0$, and $E^{\mathrm{c}}$ is a line-bundle colinear with $X$.
We refer to [19] for general properties about dynamics preserving a uniformly hyperbolic splitting. Among many interesting properties, Anosov flows always have infinitely many periodic orbits, all of them saddle type hyperbolic. The stable manifold theorem asserts that the plane bundles $E^{\mathrm{s}}\oplus E^{\mathrm{c}}$ and $E^{\mathrm{c}}\oplus E^{\mathrm{u}}$ are tangent to a pair of transverse codimension-one foliations $\mathcal{F}^{\mathrm{cs}}$ and $\mathcal{F}^{\mathrm{cu}}$ on $M$, that intersect along the $\phi$-orbits. We call them center-stable and center-unstable foliations. Every leaf of these foliations is homeomorphic to either a cylinder or a Möbius band (in case the leaf contains a periodic orbit), or to a plane.
Let $\gamma$ be a periodic orbit of an Anosov flow $\{\phi_{t}:M\to M\}_{t\in\mathbb{R}}$ and assume its center-stable/unstable leaves $\mathcal{F}^{\mathrm{cs}}(\gamma)$ and $\mathcal{F}^{\mathrm{cu}}(\gamma)$ are cylinders. In [14] Fried introduced a surgery operation, represented in Figure (1), that consists in:
@@
    Choose a blow-up $(\phi^{*},M^{*})$ of the flow $(\phi,M)$ along $\gamma$, whose restriction to the boundary $\{\phi_{t}^{*}:\partial M^{*}\to\partial M^{*}\}_{t\in\mathbb{R}}$ is Morse-Smale.
    A blow-up of the pair $(\phi,M)$ along $\gamma$ is a 3-manifold $M^{*}$ with boundary homeomorphic to a torus, equipped with a flow $\{\phi_{t}^{*}:M^{*}\to M^{*}\}_{t\in\mathbb{R}}$ and a map $\pi:M^{*}\to M$ such that: (i)- $\pi:\partial M^{*}\to\gamma$ is a $\mathbb{S}^{1}$-fibration; (ii)- $\pi:M^{*}\setminus\partial M^{*}\to M\setminus\gamma$ is a homeomorphism; and (iii)- $\pi\circ\phi_{t}^{*}=\phi_{t}\circ\pi$, $\forall\ t\in\mathbb{R}$. There always exists a blow-up where the flow on the boundary is a Morse-Smale flow with two attracting and two repelling periodic orbits.
@@
    For this, choose a foliation $\mathcal{C}$ by simple closed curves on $\partial M^{*}$ such that each leaf is transverse to the flow $\{\phi_{t}^{*}:\partial M^{*}\to\partial M^{*}\}_{t\in\mathbb{R}}$ and intersects once each periodic orbit in the boundary. Up to a time reparametrization near the boundary we may assume that $\phi^{*}$ preserves this foliation. Then, the quotient $\sigma:M^{*}\to N\coloneqq M^{*}/\mathcal{C}$, that consists in collapsing each leaf of $\mathcal{C}$ into a point, is a $3$-manifold where $\delta\coloneqq\partial M^{*}/\mathcal{C}$ is a simple closed curve, and there is an induced flow $\{\psi_{t}:N\to N\}_{t\in\mathbb{R}}$ that has $\delta$ as a saddle type periodic orbit.
Figure 1. Fried surgery $(\phi,M)\xrightarrow[]{(\gamma,1)}(\psi,N)$.
Topologically, this operation is an integral Dehn surgery on $M$ along the simple closed curve $\gamma$. Thus the homeomorphism type of $N$ is determined by the homology class in $\partial M^{*}$ of the leaves $C$ of the chosen foliation $\mathcal{C}$. In the basis of $H_{1}(\partial M^{*})$ given by the class $a$ of an oriented $\mathbb{S}^{1}$-fiber (meridian) and the class $b$ of an oriented closed orbit (longitude) we have that $C]=a+m\cdot b$ for some $m\in\mathbb{Z}$.
This surgery has the additional property that the resulting manifold is equipped with a flow $\{\psi_{t}:N\to N\}_{t\in\mathbb{R}}$, that is actually equivalent to $(\phi,M\setminus\gamma)$ in the complement of $\delta$. It is possible to check that the orbit $\delta$ is saddle type, and that the flow $\psi$ preserves a pair of globally defined center-stable and center-unstable foliations on $N$, induced from those corresponding to $\{\phi_{t}:M\to M\}_{t\in\mathbb{R}}$. However, this flow is not an Anosov flow (i.e. does not preserves a hyperbolic splitting) in a canonical way (see [25] for a deeper discussion). It is what is called a topological Anosov flow.
@@
We say that the previous operation is an integral Fried surgery of slope $m$ on the flow $(\phi,M)$ along the orbit $\gamma$. We use the notations
$(\psi,N)=\operatorname{\mathrm{Surg}}(\phi,M;(\gamma,m))$
$(\phi,M)\xrightarrow[]{(\gamma,m)}(\psi,N)$
to indicate that $(\psi,N)$ is obtained from $(\phi,M)$ by doing the specified surgery.
Observe that, since the foliation by $\phi$-orbits is unchanged in the complement of $\gamma$ under the Fried surgery, every $\phi$-periodic orbit in $M$ produces a unique $\psi$-periodic orbit in $N$ after surgery. We denote by $\Pi:\textrm{Per}(\phi,M)\to\textrm{Per}(\psi,N)$ the natural correspondence that associates periodic orbits before and after surgery.
Remark 3.
The previous surgery can be also performed if $\mathcal{F}^{\mathrm{cs}}(\gamma)$, $\mathcal{F}^{\mathrm{cu}}(\gamma)$ are Möbius bands. However, in the present note we will only work with Anosov flows having orientable center-stable/unstable foliations.
@@ 1.2. Graph of Anosov flows
We say that two flows are orbitally equivalent if there exists a homeomorphism between their supporting manifolds that sends the orbit foliation of one flow to the orbit foliation of the other, preserving the orientation of the orbits. Using the methods in [24], it is possible to check that the Fried surgery is a well-defined operation within the family of orbital equivalence classes of (topological) Anosov flows on closed 3-manifolds. This allows to define a graph structure as follows:
@@
Define the graph of Anosov flows [^1] See [18] for a similar graph construction within the family of closed 3-manifolds. to be the 1-complex $\mathbf{G}$ having:
  • Vertices $\mathbf{G}^{0}$: Orbital equivalence classes $(\phi,M)]$ of Anosov flows on 3-manifolds;
  • Edges $\mathbf{G}^{1}$: For each vertex $v$ in $\mathbf{G}^{0}$ choose a fixed representative $(\phi,M)$ and for each pair $\gamma\in\textrm{Per}(\phi,M)$ and $m\in\mathbb{Z}$ put an oriented edge
    $(\phi,M)]=v\xrightarrow[]{(\gamma,m)}w=[(\psi,N)]$
    connecting the class $v$ to the class $w$ of $(\psi,N)=\operatorname{\mathrm{Surg}}(\phi,M;(\gamma,m))$.
Paths in the directed graph $\mathbf{G}$ encode surgery operations within the family of orbital equivalence classes of Anosov flows, and hence the topological complexity of the graph $\mathbf{G}$ reflects somehow the complexity of the surgery operations on Anosov flows.
Remark 5.
Other versions of a graph encoding surgery operations between orbital equivalence classes of Anosov flows can be given, notably with smaller number of edges. For instance, we can define another graphs by: only considering edges $(\gamma,\pm 1)$ in $\mathbf{G}^{1}$ (and removing all the others); identifying edges $(\gamma_{1},m)\sim(\gamma_{2},m)$ in $\mathbf{G}^{1}$ whenever $\gamma_{1}$ and $\gamma_{2}$ are equivalent under a self-orbital equivalence of $(\phi,M)$; or by simply adding a non-oriented edge between two classes whenever there exist at least one surgery connecting them. The results presented in this note are adapted for the case of Definition 4 above.
Anosov flows can be transitive or not, their center-stable and center-unstable foliations can be orientable or not, and their supporting 3-manifold can be orientable or not. Since Fried surgeries preserve each of these properties, the graph $\mathbf{G}$ has many connected components. One fundamental open problem is the following:
Problem 6 (Christy-Fried-Ghys, see [20]- Problem 3.54).
Let $\mathbf{G}_{\circ}$ be the subgraph of $\mathbf{G}$ whose vertices represent Anosov flows on orientable closed 3-manifolds, dynamically transitive and having orientable center-stable and unstable foliations. Is the graph $\mathbf{G}_{\circ}$ connected ?
One way to investigate the topology of $\mathbf{G}_{\circ}$ (or even $\mathbf{G}$) consists in fixing one Anosov flow, fixing a set of $n>0$ periodic orbits, and looking at all possible surgeries along the chosen curves. This approach is studied by Bonatti and Iakovoglou in [7], for the case where the Anosov flow is a suspension (cf. Section 2.1) and $n=2$.
If $\gamma_{1}$, $\gamma_{2}$ are periodic orbits of a transitive Anosov flow $(\phi,M)$ and $m_{1}$, $m_{2}$ are integers, for each $i=1,2$ let $(\phi_{i},M_{i})=\operatorname{\mathrm{Surg}}(\phi,M;(\gamma_{i},m_{i}))$ be an Anosov flow obtained by the corresponding Fried surgery, and let $\Pi_{i}:\textrm{Per}(\phi,M)\to\textrm{Per}(\phi_{i},M_{i})$ be the map that identifies periodic orbits before and after surgery, as defined before. The methods at Section 3 of [24] allow to prove that
$\operatorname{\mathrm{Surg}}(\phi_{1},M_{1};(\Pi_{1}(\gamma_{2}),m_{2}))\ \text{is orbitally equivalent to}\ \operatorname{\mathrm{Surg}}(\phi_{2},M_{2};(\Pi_{2}(\gamma_{1}),m_{1})),$
and hence $\operatorname{\mathrm{Surg}}(\phi^{A},M_{A};(\gamma_{1},m_{1}),(\gamma_{2},m_{2}))$ is a well-defined class $(\psi,N)]$ of Anosov flow, independent of the order in which the surgeries along $\gamma_{1}$ and $\gamma_{2}$ are performed.
Surgeries on a pair of periodic orbits yield a diagram:
$\left[(\phi_{2},M_{2})\right]$$\left[(\phi,M)\right]$$\left[(\psi,N)\right]$$\left[(\phi_{1},M_{1})\right]$$(\gamma_{2},m_{2})$$(\gamma_{1},m_{1})$$(\Pi_{2}(\gamma_{1}),m_{1})$$(\Pi_{1}(\gamma_{2}),m_{2})$
We will use the notation
$(\phi,M)]\xrightarrow[]{(\gamma_{1},m_{1}),(\gamma_{2},m_{2})}[(\psi,N)]$
to indicate that $(\psi,N)$ is obtained, up to orbital equivalence, by doing surgeries of slopes $m_{1}$ and $m_{2}$ along the periodic orbits $\gamma_{1}$ and $\gamma_{2}$ of $(\phi,M)$, respectively. This represents the existence of two paths in $\mathbf{G}_{\circ}$ connecting the classes $(\phi,M)]$ and $(\psi,N)]$, each one having length $2$, as indicated in the diagram above.
The work in [7] shows that:
Theorem 7 (Bonatti-Iakovoglou).
Let $(\phi^{A},M_{A})$ be the suspension flow generated by a hyperbolic matrix $A\in\mathrm{SL}(2,\mathbb{Z})$ with $\operatorname{\mathrm{tr}}(A)\geq 3$. Then,
@@
    For every pair $(\gamma_{1},\gamma_{2})$ of different periodic orbits, there is at most finitely many pairs $(m_{1},m_{2})$ of integers such that Surg(ϕ^A,M_A;(γ_1,m_1),(γ_2,m_2)) is equivalent to a suspension Anosov flow (cf. [7, Proposition 9.1]).
@@
    There exist special pairs $(\gamma_{1},\gamma_{2})$ of periodic orbits where every surgery Surg(ϕ^A,M_A;(γ_1,m_1),(γ_2,m_2)) with $(m_{1},m_{2})\neq(0,0)$ produces an Anosov flow that is not equivalent to a suspension (cf. [7, Theorem 5]).
@@
Along the discussion at [7, Section 1.3] the authors claim that, for every fixed hyperbolic matrix $A\in\mathrm{SL}(2,\mathbb{Z})$, it seems plausible the existence of at most finitely many 4-tuples $(\gamma_{1},m_{1},\gamma_{2},m_{2})$ with $\gamma_{1}\neq\gamma_{2}$ and $(m_{1},m_{2})\neq(0,0)$ such that the surgery $\operatorname{\mathrm{Surg}}(\phi^{A},M_{A};(\gamma_{1},m_{1}),(\gamma_{2},m_{2}))$ produces a suspension Anosov flow.
In this note we describe a simple construction that, for any given suspension Anosov flow, allows to obtain infinitely many pairs of periodic orbits $\gamma_{1}$, $\gamma_{2}$ and integers $m$ where the surgery with data $\{(\gamma_{1},m),(\gamma_{2},-m)\}$ yields another suspension Anosov flow. In particular, this shows that the expected claim at Remark 8 does not hold. Moreover, when the generating matrix $A$ of the suspension flow is conjugated to its inverse, this construction provides infinitely many non-trivial loops of length 2 in $\mathbf{G}_{\circ}$ with base point at $(\phi^{A},M_{A})]$.
1.3. Infinitely many loops of length 2 based at a suspension
The main result of this note is the following statement: